OpenStudy (anonymous):

tell whether the given number is a solution of each equation. -4x+7=15;-2

OpenStudy (anonymous):

all you have to do is replace x with -2. Testing a solution is just replacing the variable with the given solution.

jhonyy9 (jhonyy9):

-4*(-2)+7=8+7=15 so resulted that is right so -2 is a solution of this equation